OpenStudy (anonymous):

Jenna earns $9.75 per hour working at the mall. If Jenna worked hours last week, how much did she earn before taxes were taken out? $239.85 $243.75 $234.00 $246.00

OpenStudy (black_ninja123):

How many hours did she work last week?

OpenStudy (black_ninja123):

@WildHorseLover2002

OpenStudy (anonymous):

i donno

OpenStudy (black_ninja123):

It should say in the word problem

OpenStudy (anonymous):

https://media.studyisland.com/cgi-bin/mimetex.cgi?24 \,\frac{3}{5}

OpenStudy (anonymous):

https://media.studyisland.com/cgi-bin/mimetex.cgi?24 \,\frac{3}{5}

OpenStudy (black_ninja123):

$9.75*24=234

OpenStudy (black_ninja123):

Do you get what I just did?

OpenStudy (anonymous):

First, convert the number of hours she worked to a decimal. Now, multiply her hourly rate by the number of hours she worked. Therefore, Jenna earned $239.85 before taxes were taken out.

OpenStudy (anonymous):

its A

OpenStudy (black_ninja123):

If you knew what to do than why would you post the question

OpenStudy (anonymous):

i didnt i did c and then a clicked a and it game me the way to do it

OpenStudy (black_ninja123):

Oh

OpenStudy (anonymous):

ya its study island

OpenStudy (black_ninja123):

I hate study island

OpenStudy (anonymous):

me too

OpenStudy (black_ninja123):

What other topics do you have left in study island

OpenStudy (anonymous):

45 in ela math 2 in ss

OpenStudy (anonymous):

my teacher just reset my ela god it sucks

OpenStudy (anonymous):

and the worse part about it is i only have 2 day left to do it

OpenStudy (anonymous):

plus ols/ i am homeschooled

OpenStudy (black_ninja123):

oh I might be able to help you in some of your math What grade are you in?

OpenStudy (black_ninja123):

I'm homeschooled too

OpenStudy (anonymous):

7th grade lol it sucks but it was fun learning life science but i kinda didnt want to learn reproduction/sperm/how plants reproduce it was wierd

OpenStudy (black_ninja123):

lol yea I'm in 8th

OpenStudy (anonymous):

i listen to music cool

OpenStudy (anonymous):

https://www.youtube.com/watch?v=TmkffMQ4OMc